how do you find the inverse function f^-1(x) for f(x)=2(3)^x?
Answer by nerdybill(5413)

f(x)=2(3)^x
y = 2(3)^x
.
Reverse x and y:
x = 2(3)^y
Solve for y:
x = 2(3)^y
x/2 = 3^y
log3(x/2) = y (that's log base 3)
.
Therefore,
f^-1(x) = log3(x/2)